a light ray travels from air (n-1) into a second medium at an angle of 20 degree. The light is refracted inside the second medium at an angle of 15 degree. What is the index of refraction of the second medium?


1
Expert's answer
2021-05-09T14:35:10-0400

"\frac{\sin20°}{\sin15°}=\frac{n_2}{n_1}=\frac{n_2}{1}=n_2\to n_2=\frac{\sin20°}{\sin15°}=1.32" . Answer